OpenStudy (anonymous):

how do you find the equation of a tangent line through a point outside a circle. Suppose the point is (6,-5) and the circle is (x-1)^2 + (y^2) = 4 , but i want to find a general formula. we know that two tangents can always be drawn to a circle from any point outside the circle

OpenStudy (amistre64):

so like this?

OpenStudy (anonymous):

YESSSSSSSSS!!!!!!!!!!!!!!

OpenStudy (anonymous):

in an xy plane

OpenStudy (amistre64):

since a tangent to the circle is always at 90 degrees to the radius.... this makes a right triangle.

OpenStudy (anonymous):

hey how do you draw attachments like that

OpenStudy (amistre64):

the hyp is the length of the center of the cirlce to the point and one leg is the radius

OpenStudy (amistre64):

i use paint in the windows accessories...save it to my desktop and retrieve it from there

OpenStudy (anonymous):

you can draw circles?

OpenStudy (amistre64):

yes; its the oval selection and hold down the shift key while you drag

OpenStudy (anonymous):

yes what program is it? windows paint?

OpenStudy (anonymous):

ok so we have a point, lets call it (x0,y0) outside a circle

OpenStudy (amistre64):

yes; windows generi paint program that comes in every installation of windows since 3.xx

OpenStudy (anonymous):

ok i will check that

OpenStudy (anonymous):

i want to actually find a general formula here

OpenStudy (anonymous):

ok so we have (x0,y0) outside a circle (x-h)^2 + (y-k)^2 = r^2

OpenStudy (amistre64):

the general formula is the pythag thrm... to find a vector

OpenStudy (amistre64):

find a vector that is parallel to the tangent from the point....

OpenStudy (anonymous):

you cant do this analytically

OpenStudy (amistre64):

you can ; but then convert it :)

OpenStudy (anonymous):

we have y - y0 = m(x-x0) ,

OpenStudy (amistre64):

the slope of the tangent to the circle at any given point is f'(x)

OpenStudy (anonymous):

also the line from the center of circle to (x,y) is y-k = -1/m (x-h)

OpenStudy (anonymous):

the normal i mean

OpenStudy (anonymous):

since (x,y) is unknown

OpenStudy (amistre64):

given a standard point on the circle for say (Xc,Yc)...

OpenStudy (anonymous):

oh then i messed up

OpenStudy (amistre64):

the point has to be within a certain domain for an outside point to have a tangent right?

OpenStudy (anonymous):

thats possible

OpenStudy (anonymous):

there are 3 points given

OpenStudy (anonymous):

(h,k) the center of the circle, (xo,yo) point outside circle (x1,y1) the point on the circle and (x2,y2) the second point on circle that intersects tangent

OpenStudy (anonymous):

so 4 points

OpenStudy (amistre64):

f'(c) that passes thru P(x1,y1) in the field and P1(xc,yc) on the circle

OpenStudy (amistre64):

if we view this from the origin...would it be simpler?

OpenStudy (amistre64):

the angle formed by the center to outside point forms an angle [a] with the x axis; the angle between that line and the radius to tangent form [b] right

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

like this so far

OpenStudy (amistre64):

a' is in the wrong spot lol

OpenStudy (anonymous):

so you can do this vectorially? or with derivatives?

OpenStudy (amistre64):

you can, but lets get a clear picture first and then add the details :)

OpenStudy (amistre64):

the perp lines slope we want is tan(a+b) right?

OpenStudy (anonymous):

tan a + b ?

OpenStudy (amistre64):

a = cos^-1(CP/r)

OpenStudy (amistre64):

yes; from my picture I got a + b = the slope of the line that is perp to the tangent....

OpenStudy (amistre64):

and tan(a+b) givs us that slope; so figure the angles... :)

OpenStudy (amistre64):

does this makes sense?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

we center the circle at the origin so we can examine it...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont understand the a = arcos (CP/r) business, where does that come from

OpenStudy (amistre64):

the angle made by the line from the center to the extPoint makes an angle [a] with the x axis...that angle is cos^-1[]

OpenStudy (anonymous):

ok , the angle is ?

OpenStudy (amistre64):

the angle is cos(a) = P.r/(|P||r|)...dot product of 2 vecotrs

OpenStudy (anonymous):

cos a = Ca' / ( CP)

OpenStudy (amistre64):

P being the outside point

OpenStudy (anonymous):

what is r ?

OpenStudy (amistre64):

cos(b) = T.P/(|P||r|)

OpenStudy (amistre64):

radius = r.....

OpenStudy (anonymous):

cos b =r / TP

OpenStudy (amistre64):

to find the tangent..... we use the pythag thrm; |T|=sqrt(|P|^2-r^2)

OpenStudy (anonymous):

ok but we want the tangent line equation

OpenStudy (anonymous):

|T| is infinity

OpenStudy (anonymous):

but that line segment, right comes from pythangrean

OpenStudy (anonymous):

and you were doing the inner product, im a bit rusty

OpenStudy (amistre64):

back...

OpenStudy (anonymous):

hey

OpenStudy (amistre64):

so if we know tha toutside point ot not; we can assign a vector to it ..p <x,y> we can also assign a vector to our radius at any given point [c]... r = <c,f(c)>. when f'(c) is parallel with r-p..... we have our tangent line

OpenStudy (amistre64):

the equation of a tangent line thru any point is given by: x = x0 + at y = y0 + bt z = 0....since we are in the xy plane

OpenStudy (anonymous):

oh i think i have it

OpenStudy (anonymous):

if youre given a circle and a point outside a circle

OpenStudy (anonymous):

nevermind

OpenStudy (amistre64):

...... but i gotta mind lol

OpenStudy (amistre64):

if we are examine it from the origin; like all good analysts; we move the point and the circle by -x and -y..... make sure to write that in a note so we know were to put it when were done...

OpenStudy (anonymous):

i guess im confused on a lot of things. where did you get tan (a+b)

OpenStudy (anonymous):

where did you get arcos (cP/r)

OpenStudy (amistre64):

i was rummaging thru the thoughts in my head.... some were good, others not so good :)

OpenStudy (anonymous):

where did point c come from ?

OpenStudy (anonymous):

ok youre starting over

OpenStudy (amistre64):

c = center; ..... or origin depending on where the circle was

OpenStudy (amistre64):

since we dont know the point ot tangency, i figure vectors are a safe way to find it

OpenStudy (anonymous):

yeah thats the trouble

OpenStudy (amistre64):

we know the origin; we have vectors splayed out from there. the vector for the radius is simply <c,f(c)> where c is the point of tangency right?

OpenStudy (amistre64):

p is the vector from the origin to the outside point..<x,y>

OpenStudy (amistre64):

we know that if we subtract p from r we get a new vector heading inthe direction of tangency...

OpenStudy (amistre64):

it has to match f'(c) in order to determine if 'c' is a match..

OpenStudy (anonymous):

one sec

OpenStudy (amistre64):

when we find the right [c] we have the f'(c) as the slope for our line; and we have a point to place in it to figure out the b of y=f'(c)x + b

OpenStudy (anonymous):

ok vector R - P has the same direction as the tangent line , i agree

OpenStudy (anonymous):

because P + (R-P) = R , where R is that point (c,f(c)) or the vector to the point

OpenStudy (anonymous):

so f ' (c) will , you mean the vector c ? or just

OpenStudy (amistre64):

the derivative of the function itself.

OpenStudy (anonymous):

so f ' ( < c , f(c) ) >

OpenStudy (anonymous):

oh , the derivative of the equation of the circle

OpenStudy (amistre64):

the circle is what i had in mind; but the vector might work just as good

OpenStudy (anonymous):

2x + 2y * y' = 0

OpenStudy (anonymous):

i prefer analytic, my vector skills are shaky :)

OpenStudy (anonymous):

so y ' = -x/y